A height of 0.81m of a column of liquid (x) exerts the same pressure as a heavy box of weight 68,040N the base area of the box is 10m^2 what is the density of liquid (x)


1
Expert's answer
2022-03-15T10:34:09-0400
"P_1=P_2"

"g\rho h=\frac{W}{A}"

"\rho =\frac{W}{ghA}"

"\rho =\rm\frac{68\:040\:N}{9.8\: N/kg*0.81\: m*10\: m^2}=857\:\rm kg/m^2"